# Simultaneous Reward Distillation and Preference Learning: Get You a Language Model Who Can Do Both

### v1

@@ A.1 Proof of Non-Deterministic Preference Relations with Reward Differences
where $\sigma(\cdot)$ denotes the sigmoid function, and $r(x,y_{w})$ and $r(x,y_{l})$ represent the implicit rewards for the winning and losing responses, respectively. Recall that the BT reward formulation under the broad Plackett-Luce family models (Plackett, 1975) is characterized by underspecification. As such, it does not impose restrictions on the reward function form, provided it satisfies equivalence relations, i.e., rewards are defined up to a prompt-dependent shift (Definition 1 in Rafailov et al. (2024)). Consequently, the expected reward differences still adhere to Proposition 1 without necessarily having BT-motivated DPO’s implicit reward formulation.
@@ A.2 Proof of Lemma 1a and 1 b
where the above equation holds for all $\forall(x,y,y^{\prime})\in\mathcal{D}_{\text{pref}}$, since DPO does not distinguish between the nature of the true preference relations in estimating the true preference probabilities, assuming $(y,y^{\prime})$ appear as preferred and dispreferred responses respectively for any context $x$. Since this RHS of Eq. 15 is simply the sigmoided difference of implicit rewards assigned by DPO to estimate the true preference probabilities, is is straightforward to see from Proposition 1 that the RHS i.e., $\frac{\pi_{\theta^{*}}(y)\pi_{\mathrm{ref}}(y^{\prime})}{\pi_{\theta^{*}}(y^{\prime})\pi_{\mathrm{ref}}(y)}\to 0$ and $P(y_{w}\succ y_{l}\mid x)$ $\sim\frac{1}{2}$, as per our definition of non-deterministic preferences. Similar to Azar et al. (2023)’s argument, we can argue here that in this case when true preference probabilities are $\sim\frac{1}{2}$, i.e., non-deterministic, DPO’s empirical reward difference estimates tend toward zero which leads to underfitting of the optimal policy $\pi_{\theta^{*}}$ during alignment. Indeed, in this case, the $beta$ parameter does not provide any additional regularization effect to prevent policy underfitting especially under finite data. This completes the proof of Lemma 1a. ∎
We can similarly prove Lemma 1b in the case when $|\mathcal{D}_{\text{nd}}|\ll N$ where $N$ is assumed to be finite. In this case, in a similar vein as Azar et al. (2023), there is more likelihood that DPO reward difference estimates are 1, i.e., $r^{*}(x,y_{w})-r^{*}(x,y_{l})=1$. As such, from Eq. 12, it is straightforward to see that the term
$\log\left(\frac{\pi_{\theta^{*}}(y\mid x)\pi_{\text{ref}}(y^{\prime}\mid x)}{\pi_{\theta^{*}}(y^{\prime}\mid x)\pi_{\text{ref}}(y\mid x)}\right)\to 0$
@@ A.4 Proof of Lemma 2
Let us first rewrite the e-DPO objective over the preference dataset $\mathcal{D}_{\text{pref}}$ and examine how the optimal policy $\pi_{\theta}$ behaves upon convergence of this objective. For simplicity, we only consider the point-wise reward based distillation of the e-DPO formulation and do not consider ensembles of rewards. Additionally, note that the e-DPO objective does not require preference labels and can apply to any response pair.
@@
Since $(y_{1},y_{2})$ represents any response pairs without requiring them to be preference labels, we can substitute them with $(y,y^{\prime})\in\mathcal{D}_{\text{nd}}\subset\mathcal{D}_{\text{pref}}$, where $y=y_{w}$ and $y^{\prime}=y_{l}$. Without losing any generality, we can now rewrite the above equation as,
@@
Now, recall from our proof of Lemma 1b where we show that the RHS term of the above equation $\log\left(\frac{\pi_{\theta^{*}}(y\mid x)\pi_{\text{ref}}(y^{\prime}\mid x)}{\pi_{\theta^{*}}(y^{\prime}\mid x)\pi_{\text{ref}}(y\mid x)}\right)\to 0$, which implies that $\frac{\pi_{\theta^{*}}(y\mid x)\pi_{\text{ref}}(y^{\prime}\mid x)}{\pi_{\theta^{*}}(y^{\prime}\mid x)\pi_{\text{ref}}(y\mid x)}\to\infty$. Indeed, when $|\mathcal{D}_{\text{nd}}|\ll N$ where $N$ is finite, unregularized scalar reward estimates of the true preference probabilities can in fact grow exceedingly large in the absence of any other regularization parameters, since $\beta$ by its own does not provide enough regularization as shown in our proof of Lemma 1a. Interestingly, similar arguments have also been made in previous works (Azar et al., 2023). The rest of this proof follows the same argumentation starting Eq. 19 assuming $10<\pi_{\text{ref}}(y)<$ 1.
@@
A.5 Gradient Derivation of the Focal-Softened Log-Odds Unlikelihood Loss
In this section, we derive and analyze DRDO loss gradient and offer insights into how to compared supervised alignment objectives such as DPO (Rafailov et al., 2023). Note that we do not analyze the reward distillation component here since it does not directly interact with the focal-softened contrastive log-"unlikelihood" term in training and since it is naturally convex considering its a squared term. Let us first rewrite our full DRDO loss, as:
@@ A.5 Gradient Derivation of the Focal-Softened Log-Odds Unlikelihood Loss
We can now draw some insights and direct comparisons of our approach with Direct Preference Optimization (DPO) (Rafailov et al., 2023). As in most contrastive preference learning gradient terms (Rafailov et al., 2023; Hong et al., 2024; Xu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024), the term $\frac{\nabla_{\theta}\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{w}\mid x)}$ in Eq. 34 amplifies the gradient when $\pi_{\theta}(y_{w}\mid x)$ is low, driving up the likelihood of the preferred response $y_{w}$. Similarly, $\frac{\nabla_{\theta}\pi_{\theta}(y_{l}\mid x)}{1-\pi_{\theta}(y_{l}\mid x)}$ penalizes overconfidence in incorrect completions $y_{l}$ when $p_{w}$ is low, encouraging the model to hike preferred response likelihood while discouraging dispreferred ones.
The key insight here is that the modulating term, $(1-p_{w})^{\gamma}$, strategically amplifies corrections for difficult examples where the probability $p_{w}$ of the correct (winning) response is low. Intuitively, unlike DPO’s fixed $\beta$ that is applied across the whole training datasetw, this modulating term amplifies gradient updates when preference signals are weak ($p_{w}\approx 0.5$) and tempering updates when they are strong ($p_{w}\approx 1$), thus ensuring robust learning across varying preference scenarios. Intuitively, when ($p_{w}\approx 1$), the model is already confident of its decision since $p_{w}$ remains high, indicating increased model confidence for deterministic preferences. In contrast, when $p_{w}$ is small, $(1-p_{w})$ remains near 1, and the term $(1-p_{w})^{\gamma}$ retains significant magnitude, especially for larger values of $\gamma$. This allows $\pi_{\theta}$ in DRDO to learn from both deterministic and non-deterministic preferences, effectively blending reward alignment with preference signals to guide optimization.
@@
However, as Lemma 2 suggests, e-DPO does not directly face this limitation during training, but the degeneracy manifests upon convergence. Under the same conditions, the DRDO loss still operates at a sample level because if the DRDO estimate of $p^{*}$ (via $p_{w}$) is close to its true value of $\sim\frac{1}{2}$, the modulating factor ensures that gradients do not vanish. This allows DRDO to continue learning from such samples until convergence when winning and losing probabilities are pushed further apart (Figure 2). Intuitively, the $\log\left(\frac{\pi_{\theta}(y_{w}\mid x)}{1-\pi_{\theta}(y_{l}\mid x)}\right)$ term is minimized precisely under this condition where the modulating term is also close to zero.
Appendix B Further Notes on Experimental Setup
@@
  • Pal et al. (2024) assume non-determinism of preferences to be correlated to edit-distances between pairwise-samples, but we do not make sure assumptions and consider both the true (oracle) rewards and edit-distances between pairs to verify the robustness of our method. Pal et al. (2024)’s theoretical framework brings insights on DPO’s suboptimality assumes small edit distance between pairwise samples and they empirically show this primarily for math and reasoning based tasks. In contrast, our evaluation framework is more general in the sense that we consider both the oracle reward difference as well as edit distance in addressing DPO’s limitations in learning from non-deterministic preferences and we evaluate on a more diverse set of prompts apart from math and reasoning tasks.
@@
B.1 Choice of Evaluation Datasets
As mentioned in Section 5, our experiments were designed to balance robustness and thoroughness with research budget constraints, and to account for properties of the task being evaluated relative to the data. The rationale for evaluating DRDO’s robustness to non-deterministic or ambiguous preferences is straightforward: the TL;DR dataset is annotated with human confidence labels, which enables the creation of the $\mathcal{D}_{hc,he}$ and $\mathcal{D}_{\ell c,\ell e}$ splits requires to test the different non-determinism settings.
@@ B.1 Choice of Evaluation Datasets
B.2 TL;DR Summarization Dataset Splits
@@
Table 5 shows the mean confidence and normalized edit distance statistics in the TL;DR dataset which we used to compute the deterministic and non-determinisitic splits. $\mathcal{D}_{all}$ represents the full data (Stiennon et al., 2020). $\mathcal{D}_{hc,he}$ and $\mathcal{D}_{\ell c,\ell e}$ represent subsets created by splitting at the 50th percentile of human labeler confidence and edit distance values. The range of labeler confidence values for the full training data range is $[1,9]$. Additionally, segmenting the training data based on the combination of confidence and edit distance thresholds do not make the splits roughly half of the full training data. This is because there are many samples that do not simultaneously satisfy the 50th percentile threshold for each metric. In reality, this choice is intentional to more robustly evaluate DRDO under more difficult preference data settings. Additionally, previous theoretical as well as empirical work (Pal et al., 2024) has shown that supervised methods like DPO fail to learn optimal policies when the token-level similarity is high in the preference pairs, especially in the beginning of the response. Therefore, we apply this combined thresholding for all our experiments on TL;DR summarization dataset.
Appendix C Preference Likelihood Analysis
@@ Appendix D Hyperparameters
Appendix E Win-Rate Evaluation Prompt Formats

### v2

@@ B.2 Proof of Non-Deterministic Preference Relations with Reward Differences
where $\sigma(\cdot)$ denotes the sigmoid function, and $r(x,y_{w})$ and $r(x,y_{l})$ represent the implicit rewards for the winning and losing responses, respectively. Recall that the BT reward formulation under the broad Plackett-Luce family models (Plackett, 1975) is characterized by underspecification. As such, it does not impose restrictions on the reward function form, provided it satisfies equivalence relations, i.e., rewards are defined up to a prompt-dependent shift (Definition 1 in Rafailov et al. (2024b)). Consequently, the expected reward differences still adhere to Proposition 2 without necessarily having BT-motivated DPO’s implicit reward formulation.
@@ B.3 Proof of Lemma 1a and 1b
where the above equation holds for all $\forall(x,y,y^{\prime})\in\mathcal{D}_{\text{pref}}$, since DPO does not distinguish between the nature of the true preference relations in estimating the true preference probabilities, assuming $(y,y^{\prime})$ appear as preferred and dispreferred responses respectively for any context $x$. Since this RHS of Eq. 33 is simply the sigmoided difference of implicit rewards assigned by DPO to estimate the true preference probabilities, it is straightforward to see from Proposition 2 that the RHS i.e., $\frac{\pi_{\theta^{*}}(y)\pi_{\mathrm{ref}}(y^{\prime})}{\pi_{\theta^{*}}(y^{\prime})\,\pi_{\mathrm{ref}}(y)}\to 1$ and $P(y_{w}\succ y_{l}\mid x)$ $\sim\frac{1}{2}$, as per our definition of non-deterministic preferences. Since the reference model $\pi_{\mathrm{ref}}$ is assumed to have full support over the output space ($\text{supp}(\pi_{\text{ref}})=\mathcal{Y}$) and is not updated and can be set to a uniform prior ($\pi_{\text{ref}}\sim\mathcal{U}(\mathcal{Y})$(Xu et al., 2024), without losing any generality, this implies that $\frac{\pi_{\theta^{*}}(y)}{\pi_{\theta^{*}}(y^{\prime})}$ must remain close to 1 to satisfy this constraint ($\frac{\pi_{\theta^{*}}(y)\,\pi_{\mathrm{ref}}(y^{\prime})}{\pi_{\theta^{*}}(y^{\prime})\pi_{\mathrm{ref}}(y)}\to 1$). In this case, the policy tends to underfit the preference distribution since the preference signals are weak and policy cannot distinguish between the preferred and the dispreferred response.
Similar to Azar et al. (2023)’s argument, we can argue here that in this case when true preference probabilities are $\sim\frac{1}{2}$, i.e., non-deterministic, DPO’s empirical reward difference estimates actually tend toward 1 which leads to underfitting of the optimal policy $\pi_{\theta^{*}}$ during alignment. Indeed, in this case, the $\beta$ parameter does not provide any additional regularization effect to prevent policy underfitting especially under finite data. This completes the proof of Lemma 1a. ∎
We can similarly prove Lemma 1b in the case when $|\mathcal{D}_{\text{nd}}|\ll N$ where $N$ is assumed to be finite. In this case, in a similar vein as Azar et al. (2023), there is more likelihood that DPO sigmoided reward difference estimates are 1, i.e., $r^{*}(x,y_{w})-r^{*}(x,y_{l})\to\infty$.
As such, from Eq. 30, it is straightforward to see that the DPO’s implicit reward difference tends to infinity, regardless of the strength of the $\beta$ parameter, as shown below:
$\log\left(\frac{\pi_{\theta^{*}}(y\mid x)\pi_{\text{ref}}(y^{\prime}\mid x)}{\pi_{\theta^{*}}(y^{\prime}\mid x)\pi_{\text{ref}}(y\mid x)}\right)\to\infty$
@@ B.5 Proof of Lemma 2
Let us first rewrite the e-DPO’s distillation objective (Fisch et al., 2024) over the preference dataset $\mathcal{D}_{\text{pref}}$ and examine how the optimal policy $\pi_{\theta}$ behaves upon convergence of this objective. For simplicity of analysis, we only consider the point-wise reward based distillation loss in the e-DPO formulation without [^9] Note that in our empirical experiments we use the full e-DPO objective with a set of three reward models. considering reward ensembles. Note that the e-DPO objective does not require preference labels and can apply to any response pair.
@@
Since $(y_{1},y_{2})$ represents any response pair without a preference label, we can substitute them with $(y,y^{\prime})\in\mathcal{D}_{\text{nd}}\subset\mathcal{D}_{\text{pref}}$, where $y=y_{w}$ and $y^{\prime}=y_{l}$. Without losing any generality, we can now rewrite the above equation as,
@@
Now, recall from our proof of Lemma 1b where we show that the RHS term of the above equation $\log\left(\frac{\pi_{\theta^{*}}(y\mid x)\pi_{\text{ref}}(y^{\prime}\mid x)}{\pi_{\theta^{*}}(y^{\prime}\mid x)\pi_{\text{ref}}(y\mid x)}\right)\to\infty$, which implies that $\frac{\pi_{\theta^{*}}(y\mid x)\pi_{\text{ref}}(y^{\prime}\mid x)}{\pi_{\theta^{*}}(y^{\prime}\mid x)\pi_{\text{ref}}(y\mid x)}\to\infty$. Indeed, when $|\mathcal{D}_{\text{nd}}|\ll N$ where $N$ is finite, unregularized scalar reward estimates of the true preference probabilities can in fact grow exceedingly large in the absence of any other regularization parameters, since $\beta$ by its own does not provide enough regularization as shown in our proof of Lemma 1a. Interestingly, similar arguments have also been made in previous works (Azar et al., 2023). The rest of this proof follows the same argumentation starting Eq. 37 assuming $10<\pi_{\text{ref}}(y)<$ 1.
@@
B.6 Gradient Derivation of the Focal-Softened Log-Odds Unlikelihood Loss
In this section, we derive and analyze DRDO loss gradient and offer insights into how to compared supervised alignment objectives such as DPO (Rafailov et al., 2024b). Note that we do not analyze the reward distillation component here since it does not directly interact with the focal-softened contrastive log-"unlikelihood" term in training and since it is naturally convex considering its a squared term. Let us first rewrite our full DRDO loss, as:
@@ B.6 Gradient Derivation of the Focal-Softened Log-Odds Unlikelihood Loss
We can now draw some insights and direct comparisons of our approach with Direct Preference Optimization (DPO) (Rafailov et al., 2024b). As in most contrastive preference learning gradient terms (Rafailov et al., 2024b; Hong et al., 2024; Xu et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024), the term $\frac{\nabla_{\theta}\pi_{\theta}(y_{w}\mid x)}{\pi_{\theta}(y_{w}\mid x)}$ in Eq. 52 amplifies the gradient when $\pi_{\theta}(y_{w}\mid x)$ is low, driving up the likelihood of the preferred response $y_{w}$. Similarly, $\frac{\nabla_{\theta}\pi_{\theta}(y_{l}\mid x)}{1-\pi_{\theta}(y_{l}\mid x)}$ penalizes overconfidence in incorrect completions $y_{l}$ when $p_{w}$ is low, encouraging the model to hike preferred response likelihood while discouraging dispreferred ones.
The key insight here is that the modulating term, $(1-p_{w})^{\gamma}$, strategically amplifies corrections for difficult examples where the probability $p_{w}$ of the correct (winning) response is low. Intuitively, unlike DPO’s fixed $\beta$ that is applied across the whole training dataset, this modulating term amplifies gradient updates when preference signals are weak ($p_{w}\approx 0.5$) and tempering updates when they are strong ($p_{w}\approx 1$), thus ensuring robust learning across varying preference scenarios. Intuitively, when ($p_{w}\approx 1$), the model is already confident of its decision since $p_{w}$ remains high, indicating increased model confidence for deterministic preferences. In contrast, when $p_{w}$ is small, $(1-p_{w})$ remains near 1, and the term $(1-p_{w})^{\gamma}$ retains significant magnitude, especially for larger values of $\gamma$. This allows $\pi_{\theta}$ in DRDO to learn from both deterministic and non-deterministic preferences, effectively blending reward alignment with preference signals to guide optimization.
@@
However, as Lemma 2 suggests, e-DPO does not directly face this limitation during training, but the degeneracy manifests upon convergence. Under the same conditions, the DRDO loss still operates at a sample level because if the DRDO estimate of $p^{*}$ (via $p_{w}$) is close to its true value of $\sim\frac{1}{2}$, the modulating factor ensures that gradients do not vanish. This allows DRDO to continue learning from such samples until convergence when winning and losing probabilities are pushed further apart (Fig. 4). Intuitively, the $\log\left(\frac{\pi_{\theta}(y_{w}\mid x)}{1-\pi_{\theta}(y_{l}\mid x)}\right)$ term is minimized precisely under this condition where the modulating term is also close to zero.
Finally, Table 3 provides pseudocode for the DRDO algorithm.
DRDO Algorithm
Input: Preference dataset $\mathcal{D}_{\text{pref}}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, initialized policy model with reward head $\pi_{\theta,\theta^{\prime}}\leftarrow\text{SFT}(\theta)\oplus r_{\theta^{\prime}}$.
Output: Optimized model parameters $\theta$ in policy $\pi_{\theta}$.
1. Train Oracle $r_{\phi}$ with loss $\mathcal{L}_{\mathcal{O}}(r_{\phi},\mathcal{D}_{\text{pref}})$ (see Eq. 5).
2. For $t=1,\ldots,T$:
      (a) For each $(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})$ in $\mathcal{D}_{\text{pref}}$:
          i. Compute $r^{*}_{1}=r_{\phi}(x^{(i)},y_{w}^{(i)})$ and $r^{*}_{2}=r_{\phi}(x^{(i)},y_{l}^{(i)})$.
          ii. Compute $\hat{r}_{1}=r_{\theta^{\prime}}(x^{(i)},y_{w}^{(i)})$ and $\hat{r}_{2}=r_{\theta^{\prime}}(x^{(i)},y_{l}^{(i)})$.
          iii. Compute knowledge distillation loss: $\mathcal{L}_{\mathrm{kd}}(r^{*},\pi_{\theta})=\mathbb{E}_{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\sim\mathcal{D}_{\text{pref}}}\Bigg{[}\underbrace{\left(r^{*}_{1}-r^{*}_{2}-(\hat{r}_{1}-\hat{r}_{2})\right)^{2}}_{\text{Reward Difference}}-\underbrace{\alpha(1-p_{w})^{\gamma}\log\left(\frac{\pi_{\theta}(y_{w}^{(i)}\mid x^{(i)})}{1-\pi_{\theta}(y_{l}^{(i)}\mid x^{(i)})}\right)}_{\text{Contrastive Log-"unlikelihood"}}\Bigg{]}.$
          iv. Update $\pi_{\theta,\theta^{\prime}}$ using $\mathcal{L}_{\mathrm{kd}}$.
3. Return: Aligned policy $\pi_{\theta}$.
Table 3: DRDO Algorithm steps. We start off with the preference dataset and an SFT-trained policy initialized with an additional linear head parameterized by $\theta^{\prime}$. Once our oracle is trained, we compute both estimated rewards for each response ($y$) from the initial policy ($\hat{r}$) as well as from the oracle ($r^{*}$). We then use $\mathcal{L}_{\mathrm{kd}}$ to update both $\theta$ and $\theta^{\prime}$ in $\pi_{\theta}$ resulting in our DRDO aligned policies.
Appendix C Further Notes on Experimental Setup
@@
  • Pal et al. (2024a) assume non-determinism of preferences to be correlated to edit-distances between pairwise-samples, but we do not make sure assumptions and consider both the true (oracle) rewards and edit-distances between pairs to verify the robustness of our method. Pal et al. (2024a)’s theoretical framework brings insights on DPO’s suboptimality assumes small edit distance between pairwise samples and they empirically show this primarily for math and reasoning based tasks. In contrast, our evaluation framework is more general in the sense that we consider both the oracle reward difference as well as edit distance in addressing DPO’s limitations in learning from non-deterministic preferences and we evaluate on a more diverse set of prompts apart from math and reasoning tasks.
@@
C.1 DRDO and e-DPO Specifics
C.2 Choice of Evaluation Datasets
As mentioned in Sec. 5, our experiments were designed to balance robustness and thoroughness with research budget constraints, and to account for properties of the task being evaluated relative to the data. The rationale for evaluating DRDO’s robustness to non-deterministic or ambiguous preferences is straightforward: the TL;DR dataset is annotated with human confidence labels, which enables the creation of the $\mathcal{D}_{hc,he}$ and $\mathcal{D}_{\ell c,\ell e}$ splits requires to test the different non-determinism settings.
@@ C.2 Choice of Evaluation Datasets
C.3 TL;DR Summarization Dataset Splits
@@
Table 4 shows the mean confidence and normalized edit distance statistics in the TL;DR dataset which we used to compute the deterministic and non-determinisitic splits. $\mathcal{D}_{all}$ represents the full data (Stiennon et al., 2020). $\mathcal{D}_{hc,he}$ and $\mathcal{D}_{\ell c,\ell e}$ represent subsets created by splitting at the 50th percentile of human labeler confidence and edit distance values. The range of labeler confidence values for the full training data range is $[1,9]$. Additionally, segmenting the training data based on the combination of confidence and edit distance thresholds do not make the splits roughly half of the full training data. This is because there are many samples that do not simultaneously satisfy the 50th percentile threshold for each metric. In reality, this choice is intentional to more robustly evaluate DRDO under more difficult preference data settings. Additionally, previous theoretical as well as empirical work (Pal et al., 2024a) has shown that supervised methods like DPO fail to learn optimal policies when the token-level similarity is high in the preference pairs, especially in the beginning of the response. Therefore, we apply this combined thresholding for all our experiments on TL;DR summarization dataset.
Appendix D Preference Likelihood Analysis
@@ Appendix E Hyperparameters
Appendix F Ablation Studies
F.1 Ablations on reward distillation and contrastive log-unlikelihood
For a more robust evaluation using a much more high-capacity oracle (GPT-4o), we randomly sampled 40 prompts from the CNN/Daily Mail test set and compute win-rates and reward margins of samples generated with a top-$p$ of 0.8 and $T=0.7$ for all baselines vs. DRDO policies trained on Reddit TL;DR. We use the Phi-3-Mini-4K-Instruct model for this experiment. We include the IPO baseline (Azar et al., 2024) with $\beta=0.1$ (or $\tau$ in their paper), the baselines without the distillation (shown as DRDO (-R)), and without the contrastive component (shown as DRDO (-C)). Additionally, we include the baseline where reward distillation term in DRDO is replaced by the DPO loss but keep the contrastive log-unlikelihood component (shown as DPO (+C)). Table 6 shows results of this experiment. Note that due to compute constraints, we only train these policies from the SFT checkpoint for 500 steps with an effective batch size of 128. For the reward estimates using GPT-4o, we only add one additional condition to the prompt shown in Fig. 5 to get scalar rewards (between 0 and 1). Fig. 9 provides the prompt format used for this experiment.
Comparison WR A (%) WR B (%) Reward A Reward B Margin A Margin B
DRDO vs. DRDO (-R) 85.0 15.0 0.24±0.26 0.07±0.08 0.22±0.22 0.09±0.04
DRDO vs. DRDO (-C) 90.0 10.0 0.25±0.23 0.05±0.07 0.23±0.22 0.06±0.03
DRDO vs. IPO 65.0 35.0 0.21±0.17 0.21±0.24 0.09±0.08 0.16±0.16
DRDO vs. DPO (+C) 80.0 20.0 0.18±0.16 0.14±0.15 0.11±0.08 0.22±0.21
Table 6: Full DRDO policies (denoted “A”) compared against various baselines (denoted “B”), including DRDO without reward distillation and DRDO without contrastive log-unlikelihood loss. Win-rates (WR) are computed using average of reward comparison for each sample and then averaged. Margins are computed using the difference of rewards.
We see that in all cases, full DRDO is the clear winner in terms of win rate. We also see that in this sample the reward distillation component contributes slightly more to the overall performance than the contrastive log-unlikelihood loss, but given the small sample size this is not a significant difference; DRDO’s performance can be attributed to a combination of both. This is reinforced by DRDO’s 80-20 performance against contrastive log-unlikelihood combined with DPO loss instead of DRDO reward distillation.
F.2 Extent of Out-of-distribution data
In order to comprehensively evaluate DRDO’s performance against baseline methods under increasing out-of-distribution (OOD) conditions, we randomly sample 1,000 prompts from the CNN/DailyMail dataset and segment these prompts into bins of 50 tokens based on prompt-token counts, spanning the full range of token lengths. Since the CNN/DailyMail dataset represents a previously unseen input distribution, this evaluation effectively measures the OOD generalization capabilities of the policies as prompt lengths (and corresponding news article lengths) increase. For this automatic evaluation, we use GPT-4o as a high-capacity judge, consistent with Sec. 5 (prompt used is shown in Fig. 7). For response sampling, we use top-$p$ of 0.8 and temperature of 0.7 for DRDO and all baselines. The trends in Fig. 6 suggest that as OOD composition (with prompt-token lengths as proxy) increases, on average DRDO policies tend to have relatively larger win-rates compared to shorter prompts over baselines like DPO and e-DPO. In contrast, we find that DPO and e-DPO win-rates tend to decrease with increase in prompt-lengths.
Appendix G Win-Rate Evaluation Prompt Formats